# Conditional Density Estimations from Privacy-Protected Data

### v3

@@ 4.2 Bayesian linear regression
We demonstrate our methods on a linear regression model, and compare it to existing work on DP regression analysis. We consider linear regression with $n$ subjects and $p$ predictors. Denote $x_{0}\in\mathbb{R}^{n\times p}$ as the design matrix without intercept terms, and let $x=(\textbf{1}_{n\times 1},x_{0})$ represent the design matrix with the intercept. Ordinary linear regression models assume that the outcomes $y$ satisfy $y|x_{0}\sim\mathcal{N}_{n}(x\beta,\sigma^{2}I_{n})$. Under the constraint of differential privacy, both outcomes $y$ and predictors $x_{0}$ are subject to calibrated noise. In a Bayesian setting, we model the predictors with $x_{0,i}\sim\mathcal{N}_{p}(m,\Sigma)$ for $i=1,2,\cdots,n$ independently. Our parameter of interest is $\beta$, which represents the $(p+1)$-dimensional vector of regression coefficients. Our experiments assume that $\sigma$, $m$, and $\Sigma$ are fixed, to illustrate our algorithm. In practice, one can also estimate these parameters from data.
@@
Private sufficient statistics. We achieve $\epsilon$-DP on confidential data $(x,y)$ by adding Laplace noise to sufficient statistics. Achieving privacy requires finite $\ell_{1}$ sensitivity on confidential data. As a result, before adding noise for privacy, we first need to clamp the predictors and the responses, and then normalize them to take values in $[-1,1]$. Let’s denote the clamped confidential data as $\tilde{x}$ and $\tilde{y}$, respectively. We then define the summary statistics of clamped data as $\tilde{s}:=\left(\frac{1}{n}\tilde{x}^{\top}\tilde{y},\frac{1}{n}\tilde{y}^{\top}\tilde{y},\frac{1}{n}\tilde{x}^{\top}\tilde{x}\right)$. We have refined the sensitivity analysis of Ju et al., (2022) to $\Delta(s)=\frac{1}{n}(p^{2}+3p+3)$. The privacy summary statistics $s_{\mathrm{dp}}$ is achieved by adding independent $\mathrm{Laplace}(0,\Delta_{1}(\tilde{s})/\epsilon)$ noise to each entry of $\tilde{s}.$ This mechanism satisfies $\epsilon$-DP, and it is connected to the measurement-error models in statistics (Carroll et al.,, 1995; Hu et al.,, 2022). Details of clamping and sensitivity analysis are in the supplementary materials.
@@
The cost of privacy. Although it has been a standard practice in many DP work (Bernstein and Sheldon,, 2018, 2019; Ju et al.,, 2022; Gong,, 2022) to achieve finite global sensitivity through clamping, this benefit of privacy protection comes at a cost of accuracy of the subsequent statistical analysis. Table 1 compares SPPE and SPLE against methods such as SMC-ABC, Data-augmentation MCMC [DA-MCMC,(Ju et al.,, 2022)], Gibbs-SS (Bernstein and Sheldon,, 2019), and RNPE (Ward et al.,, 2022) at privacy level $\epsilon=10$; we include the $\epsilon=3$ version of Table 1 in Table S2 of the Appendix.
Table 1: Estimated posterior mean and 95% credible intervals for the linear regression example using various methods: SMC-ABC, Data-augmentation MCMC [DA-MCMC,(Ju et al.,, 2022)], SPPE, SPLE, Gibbs-SS (Bernstein and Sheldon,, 2019), RNPE (Ward et al.,, 2022). Here privacy loss budget is set to $\epsilon=10$.
$\beta_{0}$ $\beta_{1}$ $\beta_{2}$
SMC-ABC -0.59 (-2.28, 0.93) -2.44 (-3.67, -0.38) 0.85 (-0.88, 2.77)
DA-MCMC -0.62 (-2.50, 0.99) -2.72 (-3.74, -0.96) 0.54 (-1.06, 2.46)
SPPE -0.51 (-2.25, 1.07) -2.40 (-3.61, -0.30) 0.90 (-0.89, 2.85)
SPLE -0.64 (-2.31, 0.96) -2.61 (-3.65, -0.98) 0.64 (-0.93, 2.48)
Gibbs-SS -0.46 (-2.22, 1.39) -0.50 (-2.08, 1.28) 0.41 (-1.68, 2.13)
RNPE -1.40 (-3.59, 0.94) -2.15 (-3.34, 0.51) 0.29 (-2.11, 2.97)
@@
Table 1 illustrates the biases introduced by measurement errors (privacy noise) and truncation (clamping). First of all, we highlight that it is necessary to design a valid inference procedure after DP data release, as a naive plug-in estimator (plugging in $s_{\mathrm{dp}}$ as $s$ into the conjugate posterior $\pi(\theta\mid s)$) gives the wrong posterior. In fact, the clamping step has destroyed the conjugacy properties enjoyed by the confidential data model. Second, achieving privacy protection comes at a cost of estimation accuracy: private data posterior $\pi(\theta\mid s_{\mathrm{dp}})$ is different from the confidential data posterior $\pi(\theta\mid x)$ even under a high loss budget of $\epsilon=10$ (small privacy noise) setting. With small privacy noise, data corruption primarily comes from the clamping step. Elevating this censoring bias is still an open problem in DP data analysis, with some recent attempts in Biswas et al., (2020); Evans et al., (2019); Covington et al., (2021).
To further characterize the privacy-utility trade-off, we compare the private data posterior distributions under several levels of privacy loss budget, in Fig. 5. The underlying confidential data $x$ is the same for each $\epsilon=0.1,0.3,1,3,10,30,100.$ For each privacy loss level, we simulate one private summary $s_{\mathrm{dp}}(\epsilon)$, and use SPPE and SPLE to approximate the private data posterior $\pi(\theta\mid s_{\mathrm{dp}};\epsilon).$ The two methods yield very similar posterior inferences. In fact, we use the same confidential data $x$ as Ju et al., (2022) and the posterior distributions in Fig. 5 follow a similar trend with that in Fig. 2 of Ju et al., (2022). For larger privacy loss budget (smaller noise), confidential data is mainly corrupted through clamping and the proposed methods can elevate the effect of this censoring bias, as $\mathbb{E}(\theta\mid s_{\mathrm{dp}})$ is closer to the confidential data expectation $\mathbb{E}(\theta\mid x).$ As $\epsilon$ get closer to 0 (larger noise, near perfect privacy), the posterior distribution is more dispersed than the prior distribution, because the privacy mechanism has injected too much noise into $s_{\mathrm{dp}}$.
@@ 5 Discussion
In this work, we propose three simulation-based inference methods to learn population parameters from privacy-protected data: SMC-ABC, SPPE, and SPLE. The latter two are neural density estimation methods. Leveraging normalizing flows and RQMC, SPPE aims to approximate the posterior-data posterior and SPLE approximates the posterior-data likelihood. Our work contributes to the growing literature on statistical analysis after DP. In particular, the framework can be applied to complex models, with intractable likelihood functions. The methodology can be especially useful for generative models, which have complex likelihood functions but are easy to simulate from, with the continuous-time SIR infection model being an example of it.
We hope our methods can catalyze DP-protected data sharing between data collectors and analysts. Our experiments and analysis demonstrate the necessity and feasibility of designing valid statistical inference procedures to correct for biases introduced by the privacy-protection mechanisms. We advocate for both an increase of sharing privacy-protected data by collectors and the use of valid inference procedures.
@@
We compare our methods to similar DP data analysis work that focuses on post processing of privacy-protected datasets (or their summary statistics). Compared with ABC (Gong,, 2022), SPPE and SPLE do not reject training samples and hence is more computationally efficient. Compared with DA-MCMC (Ju et al.,, 2022), our method does not require that the confidential data likelihood can be evaluated easily. Our methods require only that one can simulate from the prior distribution $\pi(\theta)$ and the confidential model $f(x\mid\theta).$ They also all scale linearly with sample size of the confidential database.
@@
We acknowledge the limitations of the present work and point out future directions. Using RQMC, our methods leverage the low-dimensionality of $s_{\mathrm{dp}}$ and the fact that popular DP mechanisms can be efficiently achieved with random number generators. In some applications, such as DP principle component analysis (Chaudhuri et al.,, 2013), the privacy mechanism is actually intractable to simulate from but its density is easy to evaluate. A novel perspective is to consider the whole process of generating $s_{\mathrm{dp}}$ given $\theta$ and skipping confidential data simulation as an intermediate step. This corresponds to the case of $M=1$ in Eq.(11), as analyzed by Bornn et al., (2017) for similar problems in the context of ABC. Since our method scales linearly in sample size, it might not be ideal for massive datasets, such as the Facebook URL dataset (Evans and King,, 2023), which concerns millions of users. It is of interest to develop methods that scale sub-linearly in sample size. Another interesting future direction is the distributed setting (Alparslan et al.,, 2023), where multiple parties hold parts of the data.
Supplementary Materials to
‘Conditional Density Estimations from Privacy-Protected Data’
S1 Derivations for SPLE and SPPE objectives
@@
$\ell_{\mathrm{PLE}}(\phi)=\mathbb{E}_{p(\theta,x)}\left[-\int_{\mathbb{S}}\eta(s_{\mathrm{dp}}\mid x)\log q_{\phi}(s_{\mathrm{dp}}\mid\theta)\mathrm{d}s_{\mathrm{dp}}\right].$ (S1)
@@
$\ell_{\mathrm{PLE}}(\phi)=\mathbb{E}_{\tilde{p}(\theta,x)}\left[-\int_{\mathbb{S}}\eta(s_{\mathrm{dp}}\mid x)\log q_{\phi}(s_{\mathrm{dp}}\mid\theta)\mathrm{d}s_{\mathrm{dp}}\right].$ (S2)
@@
$\ell_{\mathrm{PPE}}(\phi)=\mathbb{E}_{p(\theta,x)}\left[-\int_{\mathbb{S}}\eta(s_{\mathrm{dp}}\mid x)\log q_{\phi}(\theta\mid s_{\mathrm{dp}})\mathrm{d}s_{\mathrm{dp}}\right].$ (S3)
@@
$\ell_{\mathrm{PPE-A}}(\phi)=\mathbb{E}_{\tilde{p}(\theta,x)}\left[-\int_{\mathbb{S}}\eta(s_{\mathrm{dp}}\mid x)\log\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}})\mathrm{d}s_{\mathrm{dp}}\right],$ (S4)
@@
$\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}}):=q_{\phi}(\theta\mid s_{\mathrm{dp}})\frac{\tilde{p}(\theta)}{\pi(\theta)}\frac{1}{Z(s_{\mathrm{dp}},\phi)},\quad Z(s_{\mathrm{dp}},\phi)=\int_{\Theta}q_{\phi}(\theta\mid s_{\mathrm{dp}})\frac{\tilde{p}(\theta)}{\pi(\theta)}\mathrm{d}\theta.$
@@
$=\iint_{\mathbb{S}\times\Theta}p(s_{\mathrm{dp}})\tilde{\pi}(\theta\mid s_{\mathrm{dp}})\log\tilde{\pi}(\theta\mid s_{\mathrm{dp}})\mathrm{d}s_{\mathrm{dp}}\mathrm{d}\theta-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\int_{\mathbb{S}}\eta(s_{\mathrm{dp}}\mid x)\log\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}})\mathrm{d}s_{\mathrm{dp}}\right],$ (S5)
@@
Based on Prop. 1 in the work of Papamakarios and Murray, (2016), minimizing (S5) results in the convergence of $\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}})$ to $\tilde{\pi}(\theta\mid s_{\mathrm{dp}})$ and $q_{\phi}(\theta\mid s_{\mathrm{dp}})$ to $\pi(\theta\mid s_{\mathrm{dp}})$. ∎
S2 Proof of Proposition 4
@@
Consider a sequence of $L$ points in the time interval $[0,T]$, our privatized query can be $s_{\mathrm{dp}}=(s_{1},\ldots,s_{L})$ where each $s_{i}\sim\mathrm{Binomial}\left(n,\frac{I(t_{i})+m}{K+2m}\right)$ independently. The mechanism generating $s_{\mathrm{dp}}=(s_{1},\ldots,s_{L})$ satisfies $\epsilon$-DP, with $\epsilon=\frac{n}{m}L.$
@@
which gives the mechanism generating $s_{\mathrm{dp}}=(s_{1},\ldots,s_{L})$ satisfies $\epsilon$-DP, with $\epsilon=\frac{n}{m}L.$
S3 Experimental details

### v4

@@ 4.2 Bayesian linear regression
We demonstrate our methods on a linear regression model, and compare it to existing work on DP regression analysis like (Ju et al.,, 2022; Bernstein and Sheldon,, 2019). We consider linear regression with $n$ subjects and $p$ predictors. Denote $x_{0}\in\mathbb{R}^{n\times p}$ as the design matrix without intercept terms, and let $x=(\textbf{1}_{n\times 1},x_{0})$ represent the design matrix with the intercept. Ordinary linear regression models assume that the outcomes $y$ satisfy $y|x_{0}\sim\mathcal{N}_{n}(x\beta,\sigma^{2}I_{n})$. Under the constraint of differential privacy, both outcomes $y$ and predictors $x_{0}$ are subject to calibrated noise. In a Bayesian setting, we model the predictors with $x_{0,i}\sim\mathcal{N}_{p}(m,\Sigma)$ for $i=1,2,\cdots,n$ independently. Our parameter of interest is $\beta$, which represents the $(p+1)$-dimensional vector of regression coefficients. Our experiments assume that $\sigma$, $m$, and $\Sigma$ are fixed, to illustrate our algorithm. In practice, one can also estimate these parameters from data. Our setting is the same as that used in Ju et al., (2022).
Private sufficient statistics.
@@
We achieve $\epsilon$-DP on confidential data $(x,y)$ by adding Laplace noise to sufficient statistics. Achieving privacy requires finite $\ell_{1}$ sensitivity on confidential data. As a result, before adding noise for privacy, we first need to clamp the predictors and the responses, and then normalize them to take values in $[-1,1]$. Let’s denote the clamped confidential data as $\tilde{x}$ and $\tilde{y}$, respectively. We then define the summary statistics of clamped data as $\tilde{s}:=\left(\frac{1}{n}\tilde{x}^{\top}\tilde{y},\frac{1}{n}\tilde{y}^{\top}\tilde{y},\frac{1}{n}\tilde{x}^{\top}\tilde{x}\right)$. We have refined the sensitivity analysis of Ju et al., (2022) to $\Delta(s)=\frac{1}{n}(p^{2}+3p+3)$. The privacy summary statistics $s_{\mathrm{dp}}$ is achieved by adding independent $\mathrm{Laplace}(0,\Delta_{1}(\tilde{s})/\epsilon)$ noise to each entry of $\tilde{s}.$ This output perturbation mechanism satisfies $\epsilon$-DP. Details of clamping and sensitivity analysis are in the supplementary materials.
Posterior approximations.
We compare the 95% posterior credible intervals obtained by methods applicable to linear regression, including SMC-ABC, SPPE, SPLE, Data-augmentation MCMC (DA-MCMC) (Ju et al.,, 2022), Gibbs-SS (Bernstein and Sheldon,, 2019), and RNPE (Ward et al.,, 2022). See Table 23, and 4 for marginal posterior credible intervals of $\beta\mid s_{\mathrm{dp}}$ at privacy levels $\epsilon=10$, $1$ and $0.1$ respectively. Both tables are in Appendix E.2. We also use the Kolmogorov-Smirnov test to assess the similarity of empirical posterior marginal, shown in Figure 3 at a privacy level of $\epsilon=10$. Results from SPPE, SPLE, SMC-ABC, and DA-MCMC reach agreement on the posterior marginals. However, results from Gibbs-SS and RNPE are qualitatively different from the other four methods, yielding biased approximations for $\beta_{0}$ and $\beta_{1}$ respectively. Among the likelihood-free methods, SPLE attempts to approximate the likelihood function and is the most similar to DA-MCMC, our likelihood-based baseline.
Figure 3: Kolmogorov-Smirnov test statistics between approximations of posterior marginals, at $\epsilon=10$.
The cost of privacy.
@@
Although it has been a standard practice in many DP work (Bernstein and Sheldon,, 2018, 2019; Ju et al.,, 2022; Gong,, 2022) to achieve finite global sensitivity through clamping, this benefit of privacy protection comes at the cost of accuracy of the subsequent statistical analysis.
@@
Our experiments demonstrate the privacy-utility trade-off, which has also been demonstrated by related works on noise/privacy-aware inference. A naive plug-in estimator (plugging in $s_{\mathrm{dp}}$ as $s$ into the conjugate posterior $\pi(\theta\mid s)$) gives the wrong posterior; See the second rows in Table 2. Besides, achieving privacy protection comes at the cost of estimation accuracy: private data posterior $\pi(\theta\mid s_{\mathrm{dp}})$ is different from the confidential data posterior $\pi(\theta\mid x)$ even under a high loss budget of $\epsilon=10$ (small privacy noise) setting; See the first rows in Table 23, and 4. With small privacy noise, data corruption primarily comes from the clamping step. Evaluating this censoring bias is still an open problem in DP data analysis, with some recent attempts in Biswas et al., (2020); Evans et al., (2019); Covington et al., (2021).
@@ 5 Discussion
In this work, we propose three simulation-based inference methods to learn population parameters from privacy-protected data: SMC-ABC, SPPE, and SPLE. The latter two are neural density estimation methods. We have designed SPPE and SPLE to leverage state-of-the-art computational tools, such as normalizing flows and randomized quasi-Monte Carlo, to be computationally efficient for complex data models. SPPE aims to approximate the posterior-data posterior, and SPLE approximates the posterior-data marginal likelihood.
@@
We compare our methods to similar DP data analysis work that focuses on post processing of privacy-protected datasets (or their summary statistics). Compared with existing ABC-based analysis for DP data (Gong,, 2022), SPPE and SPLE do not reject training samples and hence are more computationally efficient. Compared with DA-MCMC (Ju et al.,, 2022), our method does not require that the confidential data likelihood can be evaluated easily. Our methods require only that one can simulate from the prior distribution $\pi(\theta)$ and the confidential model $f(x\mid\theta).$ They also all scale linearly with the sample size of the confidential database.
Our work contributes to the growing literature on statistical analysis with privatized data. In particular, our simulation-based inference framework can be applied to complex models with intractable likelihood functions and the resulting triply-intractable private data posterior. We hope our methods can catalyze DP-protected data sharing between data collectors and analysts. Our experiments and analysis demonstrate the necessity and feasibility of designing valid statistical inference procedures to correct for biases introduced by privacy-protection mechanisms. We advocate for increases in both sharing privacy-protected data by collectors and using valid inference procedures.
@@
We acknowledge the limitations of the present work and point out future directions. Our methods leverage the fact that popular DP mechanisms can be efficiently achieved with random number generators. In some tasks, such as DP principle component analysis (Chaudhuri et al.,, 2013), the privacy mechanism is actually intractable to simulate from but its density is easy to evaluate. Our method is not applicable to this type of DP mechanism. Additionally, since our method scales linearly in the sample size of the confidential database, it might not be ideal for massive datasets, such as the Facebook URL dataset (Evans and King,, 2023), which concerns millions of users. It is of interest to develop methods that scale sub-linearly in sample size. Finally, in Appendix D, we show the algorithms scale linearly with the dimension of the summary statistics and quadratically with the number of nodes per layer in the neural network. In our experiments, we found that using at most 8 layers and 50 nodes can lead to satisfactory performance. A precise understanding of how the convergence rate depends on sample size of the confidential data and the dimensionality of the summary statistics remains to be established and is likely to vary on a case-by-case basis for the choice neural density estimators.
Acknowledgments
We sincerely appreciate the valuable comments and suggestions from anonymous referees. The work of the third author was supported by the National Natural Science Foundation of China No.12171454 and Fundamental Research Funds for the Central Universities.
Appendix A Derivations for SPLE and SPPE objectives
@@
$\ell_{\mathrm{PLE}}(\phi)=\mathbb{E}_{\tilde{p}(\theta,x)}\left[-\int_{\mathbb{S}}\eta(s_{\mathrm{dp}}\mid x)\log q_{\phi}(s_{\mathrm{dp}}\mid\theta)\mathrm{d}s_{\mathrm{dp}}\right].$ (14)
@@
$\ell_{\mathrm{PPE-A}}(\phi)=\mathbb{E}_{\tilde{p}(\theta,x)}\left[-\int_{\mathbb{S}}\eta(s_{\mathrm{dp}}\mid x)\log\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}})\mathrm{d}s_{\mathrm{dp}}\right],$ (16)
@@
$\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}}):=q_{\phi}(\theta\mid s_{\mathrm{dp}})\frac{\tilde{p}(\theta)}{\pi(\theta)}\frac{1}{Z(s_{\mathrm{dp}},\phi)},\quad Z(s_{\mathrm{dp}},\phi)=\int_{\Theta}q_{\phi}(\theta\mid s_{\mathrm{dp}})\frac{\tilde{p}(\theta)}{\pi(\theta)}\mathrm{d}\theta.$ (17)
@@
$=\iint_{\mathbb{S}\times\Theta}p(s_{\mathrm{dp}})\tilde{\pi}(\theta\mid s_{\mathrm{dp}})\log\tilde{\pi}(\theta\mid s_{\mathrm{dp}})\mathrm{d}s_{\mathrm{dp}}\mathrm{d}\theta-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\int_{\mathbb{S}}\eta(s_{\mathrm{dp}}\mid x)\log\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}})\mathrm{d}s_{\mathrm{dp}}\right],$ (18)
@@
Based on Prop. 1 in the work of Papamakarios and Murray, (2016), minimizing Equation (18) results in the convergence of $\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}})$ to $\tilde{\pi}(\theta\mid s_{\mathrm{dp}})$ and $q_{\phi}(\theta\mid s_{\mathrm{dp}})$ to $\pi(\theta\mid s_{\mathrm{dp}})$. ∎
Appendix B Proof of Proposition 4
@@
Consider a sequence of $L$ points $\{t_{1},\cdots,t_{L}\}$ in the time interval $[0,T]$, our privatized query can be $s_{\mathrm{dp}}=(s_{1},\cdots,s_{L})$ where each $s_{i}\sim\mathrm{Binomial}\left(n,\frac{I(t_{i})+m}{K+2m}\right)$ independently. The mechanism generating $s_{\mathrm{dp}}=(s_{1},\cdots,s_{L})$ satisfies $\epsilon$-DP, with $\epsilon=\frac{n}{m}L.$
@@
which gives the mechanism generating $s_{\mathrm{dp}}=(s_{1},\cdots,s_{L})$ satisfies $\epsilon$-DP, with $\epsilon=\frac{n}{m}L.$
Appendix C Description of the SMC-ABC algorithm
Algorithm 3 presents the extension of the SMC-ABC algorithm (Sisson et al.,, 2007; Beaumont et al.,, 2009) to posterior estimation from the privatized data settings, by incorporating the private data simulator $f(x\mid\theta)$ and the privatized algorithm $\eta$.
Algorithm 3 Sequential Monte Carlo - Approximate Bayesian Computation (SMC-ABC)
Input: observed privatized summary statistics $s_{\mathrm{dp}}^{o}$, discrepancy function $\rho$, thresholds $\{\epsilon_{t}\}$, transition kernels $\{K_{t}\}$, simulator $f(x\mid\theta)$, and privatized algorithm $\eta$
for $t=1,2,\dots,T$ do
     for $i=1,2,\dots,N$ do
         if $t=1$ then
              Sample $\theta^{**}\sim\pi(\theta)$ independently
         else
              Sample $\theta^{*}$ from the previous population $\{\theta_{t-1}^{(i)}\}$ with weights $\{W_{t-1}^{(i)}\}$
              Perturb $\theta^{*}$ using transition kernel: $\theta^{**}\sim K_{t}(\theta\mid\theta^{*})$
         end if
         repeat
              Generate $x^{**}\sim f(x\mid\theta^{**})$
              Generate $s_{\mathrm{dp}}^{**}\sim\eta(s_{\mathrm{dp}}\mid x^{**})$
         until $\rho(s_{\mathrm{dp}}^{**},s_{\mathrm{dp}}^{o})\leq\epsilon_{t}$
         Set $\theta_{t}^{(i)}=\theta^{**}$
         Compute weights:
$W_{t}^{(i)}=\begin{cases}1,&\text{if }t=1\\ \frac{\pi(\theta_{t}^{(i)})}{\sum_{j=1}^{N}W_{t-1}^{(j)}K_{t}(\theta_{t}^{(i)}\mid\theta_{t-1}^{(j)})},&\text{if }t>1\end{cases}$
     end for
     Normalize the weights so that $\sum_{i=1}^{N}W_{t}^{(i)}=1$
     Compute effective sample size: $ESS=\left[\sum_{i=1}^{N}(W_{t}^{(i)})^{2}\right]^{-1}$
     if $ESS<E$ then
         Resample $\{\theta_{t}^{(i)}\}$ with weights $\{W_{t}^{(i)}\}$ to obtain a new population
         Set weights $\{W_{t}^{(i)}=1/N\}$
     end if
end for
return posterior estimation $\{\theta_{T}^{(i)},W_{T}^{(i)}\}_{i=1}^{N}$
Appendix D Computational Complexity Analysis
In this section, we analyze the computational complexity of the proposed (S)PLE and (S)PPE algorithms in detail. We begin by examining the computational cost of neural spline flows, which serve as the core density estimators in our framework.
Computational complexity for neural spline flows.
Denote the neural spline flow (Durkan et al.,, 2019) as $q_{\phi}(y\mid z)$ with trainable parameters $\phi$. Let $n_{l}$ represent the number of flow layers, and each layer containing $n_{h}$ hidden layers of $n_{u}$ units to generate the parameters of $K$ knots for monotonic spline. Assuming the input dimensions of $y$ and $z$ are $d_{y}$ and $d_{z}$, respectively. Each flow layer processes the concatenated input $(y,z)$ and outputs $(3K-1)$ parameter for constructing the monotonic spline for each dimension of $y$ (see Appendix A.1 in Durkan et al., (2019) for details).
The computational complexity of evaluating gradient on one sample, i.e., $\nabla_{\phi}q_{\phi}(y_{i}\mid z_{i})$, is then given by
$C_{\mathrm{NSF}}=\mathcal{O}\left(n_{l}\left[(d_{z}+d_{y})n_{u}+n_{h}n_{u}^{2}+Kd_{y}n_{u}\right]\right).$ (19)
Computational complexity for PLE.
To minimize the PLE loss $\ell_{\mathrm{PLE}}(\phi)$, the gradient with respect to $\phi$ is computed over mini-batches of size $B$. For each pair $(\theta_{i},x_{i})$ in a mini-batch, the integral $I(\theta_{i},x_{i})$ from Equation (12) is estimated using $M$ pseudo-samples $\{v^{(1)},\dots,v^{(M)}\}$ generated via RQMC.
Considering a neural spline flow with $n_{l}$ layers, each containing $n_{h}$ hidden layers with $n_{u}$ units, the computational cost for processing a single mini-batch is then
$C_{\mathrm{PLE}}=\mathcal{O}\left(n_{l}BM\left[(\mathrm{dim}(\theta)+\mathrm{dim}(s))n_{u}+n_{h}n_{u}^{2}+K\mathrm{dim}(s)n_{u}\right]+n\mathrm{dim}(\mathbb{X})BM\log M\right),$ (20)
where the last term represents the complexity of generating an RQMC-based estimator with $M$ samples. Specifically, the factor $M\log M$ accounts for the cost of generating an RQMC sequence, and $\mathrm{dim}(x)=n\mathrm{dim}(\mathbb{X})$ represents the dimension of the confidential data $x\in\mathbb{X}^{n}$ with sample size $n$.
Computational complexity for PPE.
For the PPE method, the roles of the parameters $\theta$ and summary statistics $s_{\mathrm{dp}}$ are reversed compared to PLE. To estimate the normalization constant $Z(s_{\mathrm{dp}},\phi)$ in Equation (17), we using the atomic proposal method (Greenberg et al.,, 2019), approximated as
$Z(s_{\mathrm{dp}},\phi)=\int_{\Theta}q_{\phi}(\theta\mid s_{\mathrm{dp}})\frac{\tilde{p}(\theta)}{\pi(\theta)}\mathrm{d}\theta\approx\frac{1}{n_{a}}\sum_{j=1}^{n_{a}}q_{\phi}(\theta_{j}\mid s_{\mathrm{dp},j})\frac{1}{\pi(\theta_{j})},$
where $n_{a}$ is the number of atoms, and $\{\theta_{j}\}_{j=1}^{n_{a}}$ are subsamples under the current mini-batch, which can be regarded as the samples from proposal $\tilde{p}(\theta)$ (see Appendix A.2 in Greenberg et al., (2019) for details). The computational complexity for PPE is then given by
$C_{\mathrm{PPE}}=\mathcal{O}\left(n_{l}BM\left[(\mathrm{dim}(\theta)+\mathrm{dim}(s))n_{u}+n_{a}n_{h}n_{u}^{2}+K\mathrm{dim}(s)n_{a}n_{u}\right]+n\mathrm{dim}(\mathbb{X})BM\log M\right).$ (21)
Both PLE and PPE scale linearly with the posterior dimension $\mathrm{dim}(\theta)$. Notably, while they also scale linearly with the confidential data sample size $n$, this scaling is decoupled from the complexity of the flow structure, which directly models the conditional density between the summary statistics $s$ (or $s_{\mathrm{dp}}$) and the model parameters $\theta$, rather than the high-dimensional confidential data itself.
Appendix E Experimental details